OpenStudy (anonymous):

Factor the ff. by finding the common minomial factor : 1. 2x-18 2. 3x-12 I NEED ALL YOUR HELP !! PLEASE HELP ME !! :''((

OpenStudy (anonymous):

2x-18 = 2x - 2*9 = 2(x-9) 3x - 12 = 3x - 3*4 = 2(x-4)

OpenStudy (anonymous):

* replace 2 with 3 in second

OpenStudy (anonymous):

hello . what is the meaning of this sign ? is it the exponent ?

OpenStudy (anonymous):

Ok... below is the answer. 2x-18 = 2x - 2*9 = 2(x-9) 3x - 12 = 3x - 3*4 = 3(x-4) "*" is for comment or PS like thing .... lol. I confused you

OpenStudy (anonymous):

1. 2 (x-9) 2. 3 (x-4)

OpenStudy (anonymous):

who are you queensen ? don't you see ? ms ramya teaching me !!

OpenStudy (anonymous):

lol im no good in teaching... :/

OpenStudy (anonymous):

thanks you very much ms. ramya

OpenStudy (anonymous):

it's alright i need help again

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Groupings of \[2x^{2}-x+2xy\]

OpenStudy (anonymous):

uhm, mathslover,

mathslover (mathslover):

\[\huge{x(2x-1+2y)}\] i took x common from the given expression ..

OpenStudy (anonymous):

how do you get that ?

mathslover (mathslover):

given : \[\huge{2x^2-x+2xy}\] take x common : \[\huge{x(2x-1+2y)}\] Did u get it now ?

OpenStudy (anonymous):

no, i need steps

mathslover (mathslover):

ok : \[\huge{2x^2-x+2xy}\] \[\huge{2(x*x)-(1*x)+2(x*y)}\] \[\huge{2x(x)-1(x)+2y(x)}\] \[\huge{x(2x-1+2y)}\] got it now ?

OpenStudy (anonymous):

yup, but in this equations, factor the ff. by finding the common monomial factor \[x^{2}-5x\]

OpenStudy (anonymous):

whats the answer, and how to get that ?

OpenStudy (anonymous):

well, what is the common factor in the equation? Hint: there are two \(x\)

mathslover (mathslover):

\[\huge{x^2-5x}\] \[\huge{(x*x)-(5*x)}\] \[\huge{x(x-5)}\]

OpenStudy (anonymous):

what is that ?

mathslover (mathslover):

process , i took x common

OpenStudy (anonymous):

aa ..

OpenStudy (anonymous):

mathslover, do all you required to give the answers or only the hints ?

OpenStudy (anonymous):

pls answer

mathslover (mathslover):

we are just said to give hints .. where r u having problem

OpenStudy (anonymous):

aa .. i need the answers because i have so many assignments .. :(

OpenStudy (anonymous):

asking for the answers is prolly going to take a long time if you get any at all... your best bet is to understand what everyone is teaching so you can carry out the process yourself... when you do the problems yourself, you'll find that this thing really isn't that tough.... there is nothing wrong with asking for help when you get stuck on a particular problem.... we'd be glad to help out in that case.

OpenStudy (anonymous):

:))

OpenStudy (anonymous):

still need help/explanation? :)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

factor the ff. by groupings abc plus 2c plus 3ab plus 6

OpenStudy (anonymous):

\[\large abc+2c+3ab+6 \] \[\large abc+3ab+2c+6 \] \[\large (abc+3ab)+(2c+6) \] \[\large (\color{red}{ab}c+3\color{red}{ab})+(\color{red}{2}c+\color{red}2\cdot 3) \] \[\large \color{red}{ab}(c+3)+\color{red}2(c+3) \] \[\large ab\color{blue}{(c+3)}+2\color{blue}{(c+3)} \] \[\large (ab+2)(c+3) \]